# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

### v2

@@ 5 Dataset Description
We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset (Rottensteiner et al., 2012), (2) the DFC2020 dataset (Robinson et al., 2021), (3) the RoadNet dataset (Liu et al., 2018), and (4) the Massachusetts Buildings dataset (Mnih, 2013). In the sequel, we share the details of each dataset utilised in this paper.
@@
1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm. This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset provides 38 patches (all of the size $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM). For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively. We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for semi-supervised learning.
@@
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset (Schmitt et al., 2019). It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data. The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated into ten coarser-grained classes.
@@
3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regionsThe annotation data includes road surfaces, road edges, and road centerlines for each image. We only used road surfaces for the segmentation task. the number of samples for training, validation and testing is 410, 45 and 387, respectively. In the semi-supervised learning approaches, only a quarter of the annotated training data is used, while the remaining three-quarters of the training data are employed as unlabeled data for training purposes.
@@
4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. This includes diverse buildings such as individual houses and garages, all of which are included in the labels. The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel. The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation. Each subset consists of images sized $1500\times 1500$ pixels. A quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are used in semi-supervised learning.
6 Experimental analysis
6.1 Implementation Details
@@
We implemented our methods by using the PyTorch framework. Following (Chen et al., 2018), we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  (McIntosh-Smith, 2014).
We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets. All these methods use DeepLabv3+ with ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network. To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50 on ImageNet.
@@
We comprehensively analyzed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score. Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
6.2 Quantitative Results and Analysis
@@
We evaluated the performance of the proposed approaches and some existing semi-supervised frameworks using the five performance metrics above. Experiments are conducted on the four previously mentioned datasets, namely Potsdam, DFC2020, RoadNet, and Massachusetts Building. The average results are presented in Table 1. The reason we first share average results across datasets is to provide a global performance for all utilised techniques, and the detailed performance comparison for all methods on each dataset will be discussed later. From the average results in Table 1, the proposed DiverseHead (DF) achieves the best overall performance in which there are 4 of 5 metrics reach the best compared to other methods, which are marked as red. DiverseModel shows the second-best performance among all methods in Table 1. In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.6% compared to another cross-model-based approach, CPS. Although the performance of DiverseHead (DT) is marginally lower than that of DiverseHead (DF) and DiverseModel, the three methods exhibit similar performance and surpass other compared methods. With these used remote sensing datasets, MT demonstrates notably inferior average performance across various metrics. CCT, GCT, and CPS exhibit comparable performance, yet the performance superiority of CPS is notably evident in its PA metric.
@@
Table 1: Average performance comparison with the state-of-the-art methods on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
@@
MT (Tarvainen and Valpola, 2017) 86.33% 74.90% 81.47% 65.97% 77.95%
CCT (Ouali et al., 2020) 87.10% 76.61% 82.29% 67.71% 79.19%
GCT (Ke et al., 2020) 87.43% 75.53% 82.04% 67.19% 78.63%
CPS (Chen et al., 2021) 88.03% 75.90% 85.17% 68.10% 80.13%
DiverseModel (Ma et al., 2023) 88.77% 79.54% 85.18% 70.92% 82.02%
DiverseHead (DT) 88.51% 78.40% 84.84% 70.07% 81.29%
DiverseHead (DF) 88.97% 78.99% 85.82% 71.14% 82.08%
It is important to highlight that the DiverseHead is a significantly lightweight semi-supervised learning architecture aimed at segmentation tasks within remote sensing imagery. Table 2 presents the required number of parameters during training for each semi-supervised learning architecture. Apart from DiverseHead, all other semi-supervised learning approaches typically involve parameters exceeding three hundred megabytes during training. Aside from DiverseModel, all other methods are based on the same segmentation network namely DeepLabv3+. DiverseModel is the largest architecture although it provides competitive performance compared to other cross-model-based methods namely MT, CCT, GCT, and CPS as shown in Table 1. It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands. In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only a few small convolutional layers, eliminating the need for multiple models during training. Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table 2), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Although just a a few parameters are required, the DiverseHead proposed method outperforms this state-of-the-art with at least 1% accuracy and 3% mIoU whilst obtaining similar performance to the companion method of DiverseModel.
Table 2: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
CPS (Chen et al., 2021) MT (Tarvainen and Valpola, 2017) CCT (Ouali et al., 2020) GCT (Ke et al., 2020)
@@
DiverseModel (Ma et al., 2023) DiverseHead(DT) DiverseHead(DF) Base
@@
Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table 3. The performance of the DiverseNet family, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across all evaluated datasets, as indicated by the results. In particular, for the Potsdam and Massachusetts Building datasets, the proposed DiverseHead attains the highest performance across 4 out of 5 segmentation metrics whilst DiverseModel emerges as the second-best method based on its results. However, in the case of the RoadNet and DFC2020 datasets, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance. The performance of DiverseHead (DT) consistently falls between that of DiverseModel and DiverseHead (DF) for most used datasets apart from DFC2020. This dataset provides very low-resolution labels, distinguishing it from other datasets. This could be one of the reasons why DiverseHead (DT) outperforms DiverseHead (DF) in DFC2020.
Table 3: Performance comparison with the state of the arts on four datasets. DT and DF present Dropout and dynamic freezing, respectively.
@@
Potsdam MT(Tarvainen and Valpola, 2017) 81.98% 73.66% 78.39% 63.07% 75.95%
CCT(Ouali et al., 2020) 82.66% 74.64% 77.62% 64.16% 76.10%
GCT(Ke et al., 2020) 83.99% 75.65% 80.81% 65.80% 78.14%
CPS(Chen et al., 2021) 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseModel(Ma et al., 2023) 85.76% 76.75% 83.45% 67.85% 79.96%
@@
DFC2020 MT(Tarvainen and Valpola, 2017) 78.64% 59.59% 73.57% 50.44% 65.85%
CCT(Ouali et al., 2020) 79.71% 59.87% 76.40% 51.04% 67.13%
GCT(Ke et al., 2020) 80.84% 61.47% 71.43% 52.17% 66.07%
CPS(Chen et al., 2021) 81.49% 61.74% 79.46% 53.20% 69.49%
DiverseModel(Ma et al., 2023) 81.87% 62.20% 80.69% 53.71% 70.25%
@@
RoadNet MT(Tarvainen and Valpola, 2017) 94.53% 86.77% 87.94% 78.81% 87.35%
CCT(Ouali et al., 2020) 95.45% 89.15% 89.91% 81.97% 89.53%
GCT(Ke et al., 2020) 95.20% 86.27% 90.93% 80.36% 88.54%
CPS(Chen et al., 2021) 95.66% 88.84% 90.95% 82.46% 89.88%
DiverseModel(Ma et al., 2023) 96.84% 93.12% 92.58% 87.11% 92.84%
DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
Massachusetts MT(Tarvainen and Valpola, 2017) 90.16% 79.58% 85.97% 71.55% 82.65%
CCT(Ouali et al., 2020) 90.59% 82.77% 85.23% 73.67% 83.98%
GCT(Ke et al., 2020) 89.67% 78.75% 85.01% 70.42% 81.76%
CPS(Chen et al., 2021) 89.95% 77.26% 87.33% 70.07% 81.98%
DiverseModel(Ma et al., 2023) 90.62% 86.07% 84.01% 75.00% 85.03%
DiverseHead (DT) 91.21% 82.73% 86.98% 74.67% 84.80%
DiverseHead (DF) 91.43% 83.00% 87.45% 75.16% 85.17%
@@
In order to prove the efficiency of the multi-head framework, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS). The segmentation framework consists of only a single head. During each training iteration, the current model’s predictions for unlabeled data are used to generate pseudo labels through an $argmax$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table 4, only trained by labelled data (part of each dataset) without using unlabelled data. The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table 4 to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the Base model, especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 6.91% higher than that of Base and 5.12% higher than that of SHS for Potsdam. While the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF), its performance is still much better than that of SHS and Base. Similarly, DiverseHead demonstrates superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where it exhibits a significant improvement of 6.22% over Base.
@@
Table 4: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
@@
RoadNet Base 94.86% 88.22% 88.31% 80.11% 88.26%
SHS 95.32% 87.08% 90.82% 80.95% 88.91%
DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
Table 5: Ablation study on the number of heads for DiverseHead with dynamic freezing
@@
RoadNet                       8 96.51% 91.36% 92.55% 85.67% 91.95%
                      10 96.68% 91.99% 92.73% 86.33% 92.36%
                      12 96.66% 92.12% 92.58% 86.31% 92.35%
Average                       8 91.22% 84.61% 88.20% 77.29% 86.35%
                      10 91.33% 85.57% 87.80% 77.98% 86.66%
                      12 91.18% 85.33% 87.57% 77.64% 86.42%
An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation. The results in Table 5 indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance though the performance difference between the versions with different numbers of heads is small on both datasets. Thus, we chose the 10-head version of DiverseHead for all experiments in this paper.
Table 6: Performance comparison of the DiverseModel with its constituent networks on Potsdam dataset

### v3

@@ V Dataset Description
We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset [56], (2) the DFC2020 dataset [57], (3) the RoadNet dataset [58], and (4) the Massachusetts Buildings dataset [59]. In the sequel, we share the details of each dataset utilised in this paper.
@@
1) ISPRS Potsdam Semantic Labelling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The true orthophoto (TOP) and DSM modalities have a ground sampling distance of 5 cm. Six land cover classes were identified by hand annotation of this dataset: impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset has 38 patches, each measuring $6000\times 6000$ pixels. The patches include orthorectified optical pictures of red, green, and blue bands, as well as infrared (IR) and matching digital surface models (DSM). All of these data tiles were divided into $512\times 512$ patches for computational reasons, yielding 3456, 201, and 1815 samples for the training, validation, and test sets, respectively. We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for all SSL approaches.
@@
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset, which is based on the SEN12MS dataset [60]. It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps with ten coarser-grained classes. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, one-fifth of the labelled data is for training, while the remaining four-fifths of the data is employed as unlabelled data.
@@
3) RoadNet is a benchmark dataset for roadnet detection with 0.21-m spatial resolution.It includes RGB images and related road surface maps for the segmentation task. The number of samples for training, validation and testing is 410, 45 and 387, respectively. A quarter of the annotated training data is used for the supervised part, while the remaining three-quarters of the training data are employed as unlabelled data for SSL training purposes.
@@
4) Massachusetts Buildings dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. It consists of 151 aerial RGB images with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel and corresponding building masks with the size of $1500\times 1500$ pixels. The dataset was split into 137 training, 10 testing, and 4 validation images with labels. One-quarter of the annotated training images are used for supervision, while the remaining three-quarters are treated as unlabelled for SSL.
VI Experimental analysis
VI-A Implementation Details
@@
Our approaches are implemented using PyTorch. Following [37], we employed a polynomial learning rate policy with a mini-batch SGD optimizer, where the current learning rate is calculated as the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  [61].
@@
For a fair comparison, both the state-of-the-art methods and the proposed DiverseHead use DeepLabv3+ with a ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network. We comprehensively analysed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score. Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
VI-B Quantitative Results and Analysis
@@
We evaluate the proposed approaches, along with classic and state-of-the-art SSL methods, using the five performance metrics outlined above, across four remote sensing imagery segmentation datasets: Potsdam, DFC2020, RoadNet, and Massachusetts Building. The average results across four datasets are presented in Table I. It presents an overall performance of the proposed methods using a single network in comparison to traditional SSL methods that employ dual networks. Since some state-of-the-art SSL methods in computer vision, such as UniMatch, iMAS, and AugSeg, are designed for RGB image segmentation and rely on RGB-based datasets for pretraining, these methods are not included in the average performance across the four datasets, including multi-band remote sensing data. However, a detailed performance comparison of these state-of-the-art and classic SSL methods across various datasets, including both multi-band and RGB remote sensing imagery, will be presented and discussed later in this section. From the average results in Table I, the proposed DiverseHead (DF) delivers the best overall performance, achieving the highest score in 4 out of 5 metrics, which are highlighted in red. DiverseModel demonstrates the second-best performance across most of the metrics. In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.48% compared to another network-perturbation-based approach, CPS. Although the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF) and DiverseModel, these three methods show similar average performance, but both of them outperform the other compared methods. With these used remote sensing datasets, MT demonstrates notably inferior average performance across various metrics. CCT, GCT, and CPS exhibit comparable performance, yet the performance superiority of CPS is notably evident in its PA metric. It is important to highlight that DiverseHead is a very lightweight method and reaches a very competitive average performance compared to others.
@@
Table I: Average performance comparison with the classic SSL methods with multiple networks on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
@@
MT [50] 86.35% 74.93% 81.53% 66.03% 77.99%
CCT [48] 87.14% 76.49% 82.50% 67.76% 79.23%
GCT [49] 87.45% 75.35% 82.31% 67.19% 78.65%
CPS [24] 88.06% 75.97% 85.27% 68.23% 80.22%
DiverseModel [25] 88.77% 79.54% 85.18% 70.92% 82.02%
DiverseHead (DT) 88.69% 78.51% 85.42% 70.63% 81.64%
DiverseHead (DF) 89.00% 79.14% 85.83% 71.28% 82.17%
Table II presents the number of parameters of SSL frameworks during training. Apart from DiverseHead, all other classic SSL approaches typically involve parameters exceeding three hundred megabytes. Among these approaches, DiverseModel is the largest architecture; however, it outperforms other network-perturbation-based methods, including MT, CCT, GCT, and CPS, as demonstrated in Table I. In contrast, the proposed DiverseHead is highly lightweight, using just a single segmentation model with multiple heads, each consisting of only 2 convolutional layers, thus eliminating the need for multiple networks during training. Specifically, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table II), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Despite requiring fewer parameters, the proposed DiverseHead method surpasses the classic SSL methods by at least 1% in accuracy and 3% in mIoU, while achieving performance comparable to the largest method, DiverseModel.
Table II: The required parameter size of each semi-supervised learning approach. Base means segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
CPS [24] MT [50] CCT [48] GCT [49]
@@
DiverseModel [25] DiverseHead (DT) DiverseHead (DF) Base
@@
Specifically, we present the results of different semi-supervised learning methods applied to the two multi-band remote sensing imagery datasets, namely Potsdam and DFC2020, in the table III. The performance of the DiverseNet, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across these evaluated datasets. In particular, for the Potsdam datasets, the proposed DiverseHead (DF) attains the highest performance across 4 out of 5 segmentation metrics, whilst DiverseModel emerges as the second-best method based on its results. However, in the case of the DFC2020 dataset, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance. Images from Potsdam consist of 4 bands, while images from DFC2020 contain 13 bands. The DiverseModel exhibits greater proficiency in processing datasets with more bands, while Diverse is more effective in handling images with fewer bands.
Table III: Performance comparison with the state-of-the-art on datasets containing images with more than three bands, namely Potsdam and DFC2020. DT and DF present Dropout and dynamic freezing, respectively.
@@
Potsdam MT[50] 81.98% 73.66% 78.39% 63.07% 75.95%
CCT[48] 82.66% 74.64% 77.62% 64.16% 76.10%
GCT[49] 83.99% 75.65% 80.81% 65.80% 78.14%
CPS[24] 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseModel[25] 85.76% 76.75% 83.45% 67.85% 79.96%
@@
DFC2020 MT[50] 78.64% 59.59% 73.57% 50.44% 65.85%
CCT[48] 79.71% 59.87% 76.40% 51.04% 67.13%
GCT[49] 80.84% 61.47% 71.43% 52.17% 66.07%
CPS[24] 81.49% 61.74% 79.46% 53.20% 69.49%
DiverseModel[25] 81.87% 62.20% 80.69% 53.71% 70.25%
@@
Table IV: Performance comparison with the state-of-the-art on RGB-band datasets, namely Roadnet and Massachusetts. DT and DF present Dropout and dynamic freezing, respectively.
Model OA UA PA mIoU $F_{1}$
RoadNet MT[50] 94.61% 86.87% 88.18% 79.04% 87.52%
CCT[48] 95.58% 88.67% 90.74% 82.18% 89.69%
GCT[49] 95.29% 85.53% 91.99% 80.35% 88.64%
CPS[24] 95.81% 89.10% 91.36% 82.96% 90.21%
FixMatch [62] 96.02% 89.91% 91.63% 83.81% 90.76%
UniMatch [63] 95.93% 90.58% 90.79% 83.71% 90.69%
ICNet [62] 97.15% 93.48% 93.52% 88.19% 93.50%
iMAS [51] 97.15% 94.15% 93.03% 88.32% 93.59%
AugSeg [52] 97.36% 94.26% 93.79% 89.06% 94.03%
DWL [47] 97.88% 94.81% 95.48% 90.96% 95.14%
[…]
DiverseHead(DT) w/ Sngl-model 96.85% 92.14% 93.31% 86.91% 92.72%
DiverseHead(DF) w/ Sngl-model 96.81% 92.60% 92.80% 86.88% 92.70%
DiverseHead(DT) w/ iMAS 97.47% 94.24% 94.20% 89.40% 94.22%
DiverseHead(DT) w/ AugSeg 97.70% 95.85% 93.96% 90.50% 94.90%
DiverseHead(DT) w/ DWL 97.96% 95.33% 95.36% 91.32% 95.35%
Massachusetts MT[50] 90.16% 79.58% 85.97% 71.55% 82.65%
CCT[48] 90.59% 82.77% 85.23% 73.67% 83.98%
GCT[49] 89.67% 78.75% 85.01% 70.42% 81.76%
CPS[24] 89.95% 77.26% 87.33% 70.07% 81.98%
FixMatch [62] 90.00% 81.34% 84.41% 72.13% 82.85%
UniMatch [63] 89.86% 79.95% 84.79% 71.28% 82.30%
ICNet [62] 92.24% 84.81% 88.64% 77.34% 86.68%
iMAS [51] 92.40% 83.89% 89.90% 77.26% 86.79%
AugSeg [52] 91.27% 80.92% 88.68% 73.98% 84.62%
DWL [47] 91.92% 88.29% 86.00% 77.98% 87.13%
[…]
DiverseHead(DT) w/ Sngl-model 91.21% 82.73% 86.98% 74.67% 84.80%
DiverseHead(DF) w/ Sngl-model 91.43% 83.00% 87.45% 75.16% 85.17%
DiverseHead(DT) w/ iMAS 92.92% 85.43% 90.27% 78.84% 87.78%
DiverseHead(DT) w/ AugSeg 93.35% 87.60% 89.93% 80.47% 88.75%
DiverseHead(DT) w/ DWL 92.57% 88.49% 87.37% 79.25% 87.93%
The proposed DiverseHead with two perturbation methods is a foundational framework, requiring fewer parameters and making it highly suitable for users with limited computational resources. Also, this proposed foundational framework can be easily combined with various teacher-student SSL methods to improve their performance. Table IV presents the results of the proposed methods and various SSL methods, including state-of-the-art methods in computer vision, which incorporate various enhancement strategies, on RGB remote sensing image segmentation datasets, specifically RoadNet and Massachusetts Building. DiverseHead demonstrates superior performance compared to other classic frameworks, such as MT, CCT, GCT, CPS, FixMatch and UniMatch, while maintaining a lower parameter requirement and quicker training process. Although some currently proposed SSL methods, iMAS, AugSeg and DWL, outperform the proposed DiverseHead, integrating these state-of-the-art methods with the proposed DiverseHead idea leads to further performance improvement and reaches the best result in Table IV.
MTCCTGCTCPSFixMatchUniMatchiMASAugSegDiverseHead005050505010010010010015015015015020020020020063.9663.9663.9663.96120.78120.78120.78120.78206.4206.4206.4206.493.893.893.893.887.8387.8387.8387.83104.08104.08104.08104.08152.18152.18152.18152.1881.7281.7281.7281.7240.6540.6540.6540.65Training Time (s)
Figure 7: Training time (in seconds) for 100 iterations across different SSL methods. DiverseHead refers to the dropout version of our proposed method.
To further support the computational efficiency of the proposed method, we measured the wall-clock training time required for 100 iterations across various state-of-the-art SSL frameworks, using identical hardware and implementation settings. As shown in Figure 7, DiverseHead (dropout) achieves the lowest training time (40.65 seconds) among all compared methods. This demonstrates that our approach introduces minimal computational overhead, providing strong empirical evidence of its lightweight nature during training.
@@
In order to further prove the efficiency of multi-head supervision, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS), which consists of only a single head. During each training iteration, the current model’s predictions for unlabelled data are used to generate pseudo labels through an $argmax$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table V, only trained by labelled data (part of each dataset) without using unlabelled data. The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table V to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the Base model, especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 6.91% higher than that of Base and 5.12% higher than that of SHS for Potsdam. While the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF), its performance is still much better than that of SHS and Base. Similarly, Both DiverseHead versions demonstrate superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where DiverseHead (DT) exhibits a significant improvement of 6.8% over the Base.
To evaluate the effectiveness of the proposed perturbation methods, DT and DF, in comparison to input perturbation, we conducted supplementary experiments involving data perturbation by introducing Gaussian noise to the input images. Consistency regularisation is achieved by enforcing the network to produce consistent predictions for the original and noisy input images, which is termed Input Perturbation in Table V. Both versions of DiverseHead exhibit better performance than Input Perturbation, whose performance is similar to SHS. To use the optimal Hyperparameter applied for input perturbation for fair comparison, hyperparameter tuning experiments are conducted on two datasets to confirm the standard deviation (SD) of Gaussian noise. Figure 8 illustrates the variation in mIoU as the SD changes. The mIoU value for both Potsdam and Roadnet peaked at an SD of 0.01. Thus, we set the SD of Gaussian noise for input perturbation 0.01 for the Potsdam and RoadNet datasets.
@@
Table V: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
@@
RoadNet Base 95.17% 87.62% 89.81% 80.73% 88.71%
SHS 95.44% 87.31% 91.17% 81.37% 89.20%
Input Perturbation 95.31% 87.42% 90.52% 81.03% 88.94%
DiverseHead (DT) 96.85% 92.14% 93.31% 86.91% 92.72%
DiverseHead (DF) 96.81% 92.60% 92.80% 86.88% 92.70%
Figure 8: Variation in mIoU as the standard deviation of the Gaussian noise added for input perturbation increases
Table VI: Ablation study on the effect of head number in DiverseHead with dynamic freezing.
@@
RoadNet 8 96.63% 92.15% 92.42% 86.21% 92.29%
10 96.81% 92.60% 92.80% 86.88% 92.70%
12 96.80% 92.59% 92.77% 86.84% 92.68%
Average 8 91.27% 85.01% 88.14% 77.56% 86.52%
10 91.39% 85.88% 87.83% 78.25% 86.83%
12 91.25% 85.57% 87.66% 77.90% 86.59%
To investigate the influence of head count in the DiverseHead framework on performance, an ablation study was conducted on the Potsdam and RoadNet datasets using the proposed DiverseHead framework with dynamic freezing perturbation. The results in Table VI show that using 10 heads for DiverseHead achieves the best average performance across most metrics, despite the small performance differences among the variants. This supports the choice of using 10 heads as an effective design decision.
To show the efficiency of DiverseModel, consisting of three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table VII. The component models UNet, SegNet, and PSPNet use unlabelled data through the way of SHS. DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network. In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model. This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase. The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
Table VII: Performance comparison of the DiverseModel with its constituent networks on the Potsdam dataset